# Poisson Hail on a Wireless Ground

### v1

@@ 3.1 Arrival Process
Figure 1: A discrete-time illustration of the dynamics on $\cal X\subset\mathbb{R}$. The hot ground (or equivalently, $\cal X$) is illustrated by the red line segment. Each rectangle represents an arrival, where the height represents the file size, and the projection of the rectangle on the ground represents the exclusion set.
3.2 Service Discipline
@@
As already explained, the service discipline is locally FCFS, i.e., the service of customer $n$ (arriving at time $t_{n}$ at $x_{n}$) starts as soon as all customers $k$ arriving before $t_{n}$ and such that $S_{n}\cap S_{k}\neq\emptyset$ have left. The attenuation function $l$, a function from $\mathbb{R}^{+}$ to $\mathbb{R}^{+}$, is assumed to be non-increasing and bounded from above. The positive constants $b,s$, and $w$ represent the bandwidth, the signal power and the additive white Gaussian noise power, respectively. With this notation, the service rate of customer $n$ at any time $t$, when it is active and still in the system, is
$b\log\left(1+\frac{s}{w+\sum_{x_{k}\in\phi_{t}\setminus\{x_{n}\}}l(||x_{k}-x_{n}||)}\right),$ (1)
with $\phi_{t}$ the point process made of all points $x_{k}$ of customers still in the system and being currently served at time $t$ (this excludes those customers which have precedence constraints w.r.t. other customers present in the system at that time).
3.3 Construction of model in $\mathbb{R}^{2}$ with Independent File Size
For this first description, we focus on the case with state space in $\mathbb{R}^{2}$, with Poisson space-time arrivals with intensity $\lambda$ and with service demands (file sizes) which are i.i.d. with distribution $F$ with positive and finite mean. The compact phase space is assumed to be $[-W,W]^{2}$. Its torus version is used. Concerning exclusion sets, a special case of interest is that where the exclusion set of a customer arriving at $x$ contains a ball centered at $x$ and with fixed radius $r>0$.
@@ 3.3 Construction of model in $\mathbb{R}^{2}$ with Independent File Size
We define a sequence of directed acyclic graphs (DAGs)  [^4] A DAG is a directed graph with no directed cycles., denoted by $\{G_{n}\}$, $n=0,1,2,\ldots,$ and an increasing sequence of times $\{j_{n}\}$ by induction w.r.t. $n$. The continuous-time process that we build is piecewise constant with jump times $\{j_{n}\}$, where $j_{n}$ is a strictly increasing sequence. DAG $G_{n}$ is the state DAG on the time interval $[j_{n},j_{n+1})$. A DAG is possibly non-connected. DAG $G_{n}$ features a set of vertices (customers) $V_{n}$ with acyclic directed edges $E_{n}$ between them. The service (or active) set $U_{n}$ of $G_{n}$ is the subset of vertices of $V_{n}$ without predecessor [^5] Vertex $v$ is a predecessor of $u$ if there exists a directed edge from $v$ to $u$. . Each node $v\in V_{n}$ has a location $x(v)\in[-W,W]^{2}$, an exclusion set $S(v)$, a birth time $b(v)$, and a death time $d(v)$. The vertices of $G_{n}$ have each an additional state variable which is its residual service demand (file size) at time $j_{n}$, denoted by $z_{n}(v)\in{\mathbb{R}}^{+}$.
@@
  • an i.i.d. sequence of exponential random variable $\{\widetilde{\tau}_{n}\}$ of parameter $\lambda 4W^{2}$; $\widetilde{\tau}$ represents the time that elapses between $j_{n}$ and the next potential arrival after time $j_{n}$.
  • an i.i.d. sequence of uniform random variables $\{\widetilde{X}_{n}\}$ on $[-W,W]^{2}$; $\widetilde{X}_{n}$ represents the locus of the potential next arrival after time $j_{n}$.
  • an i.i.d. sequence of random variables $\{\widetilde{S}_{n}\}$; $\widetilde{S}_{n}$ represents the footprint of the next potential arrival after time $j_{n}$.
  • an i.i.d. sequence of random variables $\{\widetilde{h}_{n}\}$ of distribution $F$; $\widetilde{h}_{n}$ represents the file size of the potential next arrival after time $j_{n}$.
The word “potential” here means that this variable can be used or discarded depending on conditions spelled out in the induction. The tilda is used to distinguish it from the effective variable. The induction is as follows. When the DAG at $j_{n}$ is $G_{n}$, define the potential departure time of node $v\in U_{n}$ as
@@
$I_{n}(v)=\sum_{u\in\Psi_{n}\setminus\{v\}}l(||v-u||).$ (4)
Let also
$q_{n}:=\min_{v\in U_{n}}p_{n}(v).$ (5)
@@
    • $V_{n+1}:=V_{n}\cup\{v^{\prime}\}$ where the extra node $v^{\prime}$ is a new arrival which is allocated location $\widetilde{X}_{n+1}$, exclusion set $\widetilde{S}_{n+1}$, and residual service time $\widetilde{h}_{n+1}$, i.e., $x(v^{\prime})=\widetilde{X}_{n+1},~{}b(v^{\prime})=j_{n+1},~{}S(v^{\prime})=\widetilde{S}_{n+1},~{}z(v^{\prime})=\widetilde{h}_{n+1}$.
    • $E_{n+1}$ is the union of $E_{n}$ and of the set of edges from ${v^{\prime}}$ to $V_{n}$, defined as follows: there is an edge from $v\in V_{n}$ to ${v^{\prime}}$ if $S(v^{\prime})\cap S(v)\neq\emptyset$. This defines $G_{n+1}$.
    • For all $v\in U_{n}$, $x_{v}$ is unchanged and the residual file size is updated to $z_{n+1}(v)=z_{n}(v)-\widetilde{\tau}_{n}r_{n}(v)$.
@@
This induction step defines $j_{n+1}$, $G_{n+1}$ and the residual service times at time $j_{n+1}$. The initial DAG $G_{0}$ is, for instance, the empty DAG and $j_{0}=0$. Up to ties, which happen with 0 probability, the above dynamics defines a piecewise constant DAG state process up to time $j=\lim_{n\to\infty}j_{n}$. It should be clear that $j=\infty$ a.s. This follows from the fact that $q_{n}$, defined above, is bounded from below by a random variable $Q_{n}$ such that the sequence $\{Q_{n}\}$ is i.i.d. and such that $\mathbb{E}Q_{n}>0$. The latter fact follows from the property that $r_{n}\leq c$ where $c$ is the constant
$c:=b\log(1+\frac{s}{w}).$
@@
Note that each DAG can be represented as a marked counting measure where the atoms of the counting measure are the locations of the customers present in the system and the marks are the precedence edges and the residual service times of these customers (all those having a predecessor in the DAG have their residual file size equal to their full initial file size).
Remark 1.
@@
In the special case where all exclusion sets contain a ball of fixed radius $R>0$, if no two customers with intersecting exclusion sets are served at the same time, the point process of active customers is a hardcore point process with balls of radius $R>0$. This implies that the total number of customers being served at any given time is bounded above by a constant that depends on $W$ and $R$. Since the function $l$ is bounded above, this shows that the interference is bounded above by a constant at any given time and locus. Hence the Shannon rate for any active customer is bounded below by a positive constant at any given time [9].
In the general case, one can still give a positive constant lower bound to the sum of the departure rates of all customers present in the system at any time, given that there is at least one customer of course:
Lemma 2.
Under the foregoing assumptions, if there is at least one customer in the system, then the sum rate of all customers in the system is bounded below by a positive constant.
Proof.
Assuming that there are $N>0$ customers in the system, the sum rate is
$\sum_{n=1}^{N}b\log\left(1+\frac{s}{w+\sum_{x_{k}\in\phi_{t}\setminus\{x_{n}\}}l(||x_{k}-x_{n}||)}\right)$
$\geq$ $b\log\left(1+\sum_{n=1}^{N}\frac{s}{w+\sum_{x_{k}\in\phi_{t}\setminus\{x_{n}\}}l(||x_{k}-x_{n}||)}\right)$
$\geq$ $b\log\left(1+\frac{Ns}{w+(N-1)L}\right),$
where $L$ is the upper bound on $l$. The fact that the function $N\to\frac{Ns}{w+(N-1)L}$ is uniformly bounded below concludes the proof. ∎
3.4 Case with Renewal Arrivals
@@
The previous construction leverages the memoryless property of the exponential distribution. It can easily be extended to the case where the arrival process is a renewal process without multiplicity (rather than a Poisson point process) where the inter-arrival times $t_{i+1}-t_{i}$ are assumed to be i.i.d. and strictly positive with a finite mean, by adding in the induction a supplementary state variable for the remaining time to the next arrival. All steps and results of the last subsection then hold for this case.
4 Stability
This section adopts the assumptions of Section 3. It makes further assumptions on the structure of the exclusion sets.
4.1 Main assumptions on exclusion sets
We present here our set of assumptions.
@@ 4.1 Main assumptions on exclusion sets
Let $k\geq 1$ and let $\{B_{i}\}_{i=1}^{k}$ be a finite collection of subsets of ${\cal X}$. Let $B_{0}=\emptyset$. There exists a positive integer $k$ and a sequence of non-empty deterministic measurable sets $B_{1},B_{2},\ldots,B_{k}$ in the space ${\cal X}$ such that
@@
  • (d) ${\mathbb{P}}(B_{i}\subseteq S)>0$, for all $i=1,2,\ldots,k$.
@@
In the special case of $S=\{x\}$ a.s. for an arrival at $x$, the assumptions cannot be jointly satisfied for $|\cal X|$$\neq 0$.
@@
In the special case where every exclusion set contains a ball of fixed radius $R>0$ centered on the arrival point, it is easy to see that the last property holds for an integer $k$ large enough, which is linked to the number of balls of radius $R$ needed to completely cover the window of interest $\cal X$. However, there may exist a collection of subsets that satisfy the same property for a smaller $k$. For example, the above assumptions hold for $k=1$ when there is a positive probability that an arrival comes with $S=\cal X$. This is the case for i.i.d. exclusion radius with exponential distribution.
Figure 2: An illustration of an underlying deterministic sequence $\{B_{i}\}_{i=1}^{k}$ with the corresponding zigzag and block structure for $\cal X\subset\mathbb{R}$. For simplicity, the arrivals are not illustrated, and here the height of $\{B_{i}\}_{i=1}^{k}$ has no physical meaning.
4.2 Block Representation
@@
In what follows, we will call $C_{n}$ a zigzag event, with $C_{n,1}$ as the “zig” event and $C_{n,2}$ the “zag’ event. We say that $2k$ is the “length” of the zigzag event (i.e. the number of the $A$ events involved), with the lengths of the zig and the zag events to be equal to $k$.
We introduce now some splitting of customers into ”blocks”, or ”cycles”.
Let $n_{0}=0$, $n_{1}=2k\cdot\min\{r\geq 1:{\mathbf{I}}(C_{2kr})=1\}$ and, for $m=1,2,\ldots$,
$n_{m+1}=n_{m}+2k\cdot\min\{r\geq 1:{\mathbf{I}}(C_{n_{m}+2kr})=1\}.$ (7)
Let $T_{0}=0$ and $T_{m}:=t_{n_{m}}$ be the corresponding thinning of the arrival process. For $m\geq 0$, let
$Y_{m}:=\{((x_{n},t_{n}-t_{n_{m}}),(S_{n},h_{n})),n=n_{m}+1,n_{m}+1,n_{m}+2,\ldots,n_{m+1}\}.$
@@
  • The last customer of this block (which also has an exclusion set containing $B_{k}$) is the closing customer of block $m$; it arrives at $\hat{T}_{m}:=t_{n_{m+1}-1}$.
Note that these blocks have independent but not identically distributed components. For example, the $k$ first arrivals of a block have exclusion sets $S_{1},\ldots,S_{k}$ such that $B_{k-i+1}\subseteq S_{i}$ for all $i=1,\ldots,k$.
Fig. 2 illustrates a zigzag and block structure based on a deterministic set of $\{B_{i}\}_{i=1}^{k}$ on $\cal X\subset\mathbb{R}$. The sequence of arrivals has exclusion sets that cover the sequence $\{B_{i}\}_{i=1}^{k}$ and are omitted. Such sets can easily be constructed for systems in $\mathbb{R}^{d}.$ One can see that the blocks are formed by “zag-zig” events and not by “zig-zag” events.
The proof of the next proposition follows directly from the strong Markov property.
@@
The sequence of random elements $\{Y_{m}\}_{m\geq 1}$ is i.i.d. and does not depend on $Y_{0}$.
@@
As already mentioned, there are several other options for defining zigzag events while preserving the properties identified above and below. One of them is that with no-doubling the last event in the zig. For example, if $k=1$ (i.e., we have customers with exclusion set equal to the whole space), then we do not have to double this event and all the properties identified still hold. The specific structure proposed here is retained for its simplicity and symmetry. This concept comes from tropical mathematics and is linked to what is now called the tropical matrix rank. See, e.g., [10].
Note also that, in Subsection 4.1, we may add for free a further assumption (e) to assumptions (a)–(d):
@@
Indeed if (e) fails for some  $i\neq j$, then we may merge these two sets and consider a new collection of $(k-1)$ events $B$’s instead.
There are also other options for defining the blocks. For example, we may deal with smaller blocks, by defining $\{n_{m}\}$ inductively as
@@
However, we prefer the definition (7) because the variant with (8) leads to a more complicated construction.
The index of the next block starts as early as $n_{m}+2k$.
4.3 Reduction to a Lindley type recurrence relation
In view of the service discipline and the definition of blocks, customers are served block by block, i.e., the service of the customers of block $m+1$ may start only when the service of the closing customer of block $m$ is completed, and at that time, all customers of block $m$ have already left. This can be exploited to derive a Lindley-type evolution equation.

### v2

@@ II-A Arrival Process
Figure 1: An illustration of the dynamics on $\cal X\subset\mathbb{R}$. The hot ground (or equivalently, $\cal X$) is illustrated by the red line segment. Each rectangle represents an arrival, where the height represents the (residual) file size, and the projection of the rectangle on the ground represents the exclusion set.
II-B Service Discipline and Service Rate
The service discipline is locally FCFS: the service of customer $n$ (arriving at time $t_{n}$ at $x_{n}$) starts as soon as all customers $k$ arriving before $t_{n}$ and such that $S_{n}\cap S_{k}\neq\emptyset$ have left.
@@ II-B Service Discipline and Service Rate
Throughout the paper, an attenuation function $l$ is assumed to be given. This is a function from $\mathbb{R}^{+}$ to $\mathbb{R}^{+}$, which is assumed to be non-increasing and bounded from above. The positive constants $b,s$, and $w$ represent the bandwidth, the signal power and the additive white Gaussian noise power, respectively. With this notation, the service rate of customer $n$ at any time $t$, when it is active and still in the system, is the Shannon rate
[…]
with $\phi_{t}$ the point process made of all points $x_{k}$ of customers still in the system and being currently served at time $t$. $\|\cdot\|$ denotes the Euclidean metric of $\cdot\in\mathbb{R}^{d}$. We make two observations in the following regarding the service rate.
II-B1 Upper Bound on Service Rate
Shannon rates are bounded from above in the sense that for all spatial configurations, the service rate of any customer is upper bounded by the constant $c:=b\log\left(1+s/w\right)$, which is the rate in the absence of interference.
II-B2 Lower Bound on Sum Service Rate
One can give a positive constant lower bound to the sum of the departure rates of all customers present in the system at any time, given that there is at least one customer, as shown in the next lemma.
Lemma 1.
Under the monotonicity and boundedness assumptions on the attenuation function, if there is at least one active customer in the system, then the sum rate of all active customers in the system is bounded below by a positive constant.
Proof.
Assuming that there are $N>0$ active customers in the system, the sum rate is
[…]
$\stackrel{{\scriptstyle\text{\scriptsize(b)}}}{{\geq}}$ $b\log\left(1+\frac{Ns}{w+(N-1)L}\right),$
where $L$ is an upper bound on $l$. Step (a) follows from the inequality $\prod_{n=1}^{N}(1+a_{n})\geq 1+\sum_{n=1}^{N}a_{n}$ for $a_{n}\geq 0$. Step (b) follows from the upper bound on $l$. The fact that the function $N\to{Ns}/{(w+(N-1)L)}$ is uniformly bounded below concludes the proof. ∎
II-C Construction of the Model
@@
We construct the model for ${\cal X}\subset\mathbb{R}^{d}$, with Poisson arrivals with intensity $\Lambda$. We assume here that $\nu$ has its first marginal $F$ with finite mean. So, the service demands (file sizes) are i.i.d. with distribution $F$. We define a sequence of directed acyclic graphs (DAGs)  [^1] A DAG is a directed graph with no directed cycles., denoted by $\{G_{n}\}$, $n=0,1,2,\ldots,$ and an increasing sequence of times $\{j_{n}\}$ by induction w.r.t. $n$. The continuous-time process that we build is piecewise constant with jump times $\{j_{n}\}$, where $j_{n}$ is a strictly increasing sequence. DAG $G_{n}$ is the state DAG on the time interval $j_{n},j_{n+1})$. A DAG is possibly non-connected. DAG $G_{n}$ has a set of vertices (customers) $V_{n}$ with acyclic directed edges $E_{n}$ between them. The service (or active) set $U_{n}$ of $G_{n}$ is the subset of vertices of $V_{n}$ without predecessor [^2] Vertex $v$ is a predecessor of $u$ if there exists a directed edge from $v$ to $u$. . Each node $v\in V_{n}$ has a location $x(v)\in{\cal X}$, an exclusion set $S(v)$, a birth time $b(v)$, and a death time $d(v)$. The vertices of $G_{n}$ have each an additional state variable which is its residual service demand (file size) at time $j_{n}$, denoted by $z_{n}(v)\in{\mathbb{R}}^{+}$.
@@
  • The sequence of i.i.d. exponential random variables $\{\widetilde{\tau}_{n}\}$ of parameter $\Lambda$; $\widetilde{\tau}_{n}$ represents the time that elapses between $j_{n}$ and the next potential arrival after time $j_{n}$ (by “potential”, we mean the following: if $\widetilde{\tau}_{n}\geq q_{n}$ defined below, then $\widetilde{\tau}_{n}$ is not used and at time $j_{n+1}$, a new exponential random variable $\widetilde{\tau}_{n+1}$ is sampled to locate the next potential arrival epoch seen from time $j_{n+1}$. This representation is fine because of the memoryless property of exponential random variables. If $\widetilde{\tau}_{n}<q_{n}$, then the next arrival takes place at time $j_{n}+\widetilde{\tau}_{n}$ and then, the random variable $\widetilde{\tau}_{n}$ is used “effectively”).
  • an i.i.d. sequence of uniform random variables $\{\widetilde{X}_{n}\}$ on $\cal{X}$; $\widetilde{X}_{n+1}$ represents the locus of the potential next arrival after time $j_{n}$.
  • an i.i.d. sequence of pairs of random variables $\{(\widetilde{R}_{n},\widetilde{h}_{n})\}$, where $\widetilde{S}_{n+1}=B(\widetilde{X}_{n+1},\widetilde{R}_{n+1})$ represents the footprint of the potential next arrival after time $j_{n}$, and $\widetilde{h}_{n+1}$ its file size.
All the sequences of random variables above are mutually independent, while dependence between $\widetilde{R}_{i}$ and $\widetilde{h}_{i}$ is allowed.
The word “potential” here means that this variable can be used or discarded depending on the conditions spelled out in the induction. The tilda is used to distinguish it from the effective variable. The induction is as follows. When the DAG at $j_{n}$ is $G_{n}$, define the potential departure time of node $v\in U_{n}$ as
@@
[…]
the total interference power seen at $v$. Note that $\Psi_{n}=\phi_{j_{n}}$ (previously defined in Eq (2)). Let
$q_{n}:=\min_{v\in U_{n}}p_{n}(v)$ (6)
denote the time elapsed until the next departure given the DAG at $j_{n}$.
@@
    • $V_{n+1}:=V_{n}\cup\{v^{\prime}\}$ where the extra node $v^{\prime}$ is a new arrival, with an assigned location $\widetilde{X}_{n+1}$, an exclusion set $\widetilde{S}_{n+1}$, and residual service time $\widetilde{h}_{n+1}$, i.e., $x(v^{\prime})=\widetilde{X}_{n+1},~b(v^{\prime})=j_{n+1},~S(v^{\prime})=\widetilde{S}_{n+1},~z(v^{\prime})=\widetilde{h}_{n+1}$.
    • $E_{n+1}$ is the union of $E_{n}$ and the set of edges from $V_{n}$ to $v^{\prime}$}, defined as follows: there is an edge from $v\in V_{n}$ to $v^{\prime}$} if $S(v^{\prime})\cap S(v)\neq\emptyset$ (to enforce the local FCFS policy). This defines $G_{n+1}$.
    • For all $v\in U_{n}$, $x(v)$ is unchanged and the residual file size is updated to $z_{n+1}(v)=z_{n}(v)-\widetilde{\tau}_{n}r_{n}(v)$.
@@
This induction step defines $j_{n+1}$, $G_{n+1}$ and the residual service times at time $j_{n+1}$. The initial DAG $G_{0}$ is, for instance, the empty DAG and $j_{0}=0$. Up to ties, which occur with 0 probability, the above dynamics defines a piecewise constant DAG state process up to time $j=\lim_{n\to\infty}j_{n}$.
Some basic observations are in order with respect to the model and its construction.
  1. 1.
    If the initial condition is a finite DAG, then $j=\infty$ a.s. In the empty graph initial condition case, this follows from the fact that for any $k>0$, at the arrival time $t_{k}$ of the $k$-th customer, at most $k$ departures can have taken place. Since $t_{k}$ tends to infinity a.s. as $k\to\infty$, we obtain that no finite accumulation points can exist for the sequence $\{j_{n}\}_{n}$ a.s. This can directly be extended to the case where the initial condition is any finite DAG.
  2. 2.
@@
    Each DAG can be represented as a marked counting measure where the atoms of the counting measure are the locations of the customers present in the system and the marks are the precedence edges and the residual service times of these customers (all those having a predecessor in the DAG have their residual file size equal to their full initial file size).
  3. 3.
@@
    In the special case where ${\cal X}=[-Q,Q]^{d}$ and all exclusion sets contain a ball of fixed radius $\overline{R}>0$, no two customers with intersecting exclusion sets are served at the same time, so that the point process of active customers (at any time) is a hardcore point process with balls of radius $\overline{R}>0$. This implies that the total number of customers that are served at any given time is bounded above by a constant that depends on $Q$ and $\overline{R}$. This has several implications. First, even if the initial DAG is infinite, we still have $j=\infty$ a.s. in the last construction algorithm. Indeed, we have at most a certain number of customers being served without interruption of service, each with a rate bounded from above by $c$ (see the previous subsection), and it is easy to see that this implies that there can be no finite accumulation point of departures a.s. under the foregoing Poisson assumptions. Second, since the function $l$ is bounded above, this also implies that the interference is bounded above by a constant at any given time and locus. Hence, the Shannon rate for any active customer is also bounded from below by a positive constant at any given time [9].
II-D Generalizations
@@
The previous construction leverages the memoryless property of the exponential distribution of the interarrival time. It can easily be extended to the case where the arrival process is a renewal process without multiplicity (rather than a Poisson point process) where the inter-arrival times $t_{i+1}-t_{i}$ are assumed to be i.i.d. and strictly positive with a finite mean, by adding in the induction a supplementary state variable for the remaining time to the next arrival. All the steps and results of the last subsection then hold for this case. It is also easy to extend this framework to the case where footprints are random compact sets that contain balls.
III Stability
The aim of this section is to establish the main stability result of the paper. This is done by reduction of the dynamics described in the last sections to a Lindley type recurrence equation. This reduction relies on assumptions on the arrival process and the exclusion sets which we call zigzag events. Under this assumption, the arrival process can be decomposed into i.i.d. blocks such that the waiting time of the first customer in a block can be determined from that of the last customer in the preceding block. This leads to the identification of an embedded Markov chain on the positive real line, whose positive recurrence can be assessed in a direct way by some natural Lyapunov function argument. The section adopts the assumptions of Section II.A and Section II.B. It makes further assumptions on the structure of the exclusion sets.
III-A Main Statistical Assumptions
III-A1 Assumptions on the Arrival Process
We recall here for easy access the following default assumptions for this section:
Assumption A1. The arrival process $\sum\delta_{t_{i}}$ is Poisson with intensity $\Lambda$ and i.i.d. marks with probability distribution $\mu\times\nu$ on ${\mathcal{X}}\times\mathbb{R}^{2}$.
Assumption A2. The marks $(R_{i},h_{i})$ (which are i.i.d.) are such that the $h_{i}$ random variables have finite mean.
III-A2 Assumptions on Exclusion Sets
We present here our third set of assumptions:
@@ III-A2 Assumptions on Exclusion Sets
Assumption A3. There exists a positive integer $k$ and a sequence of non-empty deterministic measurable sets $B_{1},B_{2},\ldots,B_{k}$ in the space $\cal X$} such that
@@
  • (d) ${\mathbb{P}}(B_{i}\subseteq S)>0$, for all $i=1,2,\ldots,k$, where $S=B(x,R)$ (see Subsection II-A).
@@
Note that conditions (a)-(c) are always satisfied (for instance, with $k=1$ and $B_{1}={\cal X}$), so that the key ingredient in this assumption is (d) in fact. For example, in the special case of $S=\{x\}$ a.s. for an arrival at $x$, the assumptions cannot be jointly satisfied for $|{\mathcal{X}}|\neq 0$.
@@
In the special case where every exclusion set contains a ball of fixed radius $\overline{R}>0$ centered on the arrival point, it is easy to see that the last property holds for an integer $k$ large enough, which is linked to the number of balls of radius $\overline{R}$ needed to completely cover the window of interest $\cal X$. However, there may exist a collection of subsets that satisfy the same property for a smaller $k$. For example, the last assumptions hold for $k=1$ when there is a positive probability that an arrival comes with $S=\cal X$. This is the case for i.i.d. exclusion radii with exponential distribution.
Figure 2: An illustration of an underlying deterministic sequence $\{B_{i}\}_{i=1}^{k}$ with the corresponding zigzag structure for $\cal X\subset\mathbb{R}$.
For all the following results, Assumptions 1-3 are assumed to hold.
III-B Block Representation
@@
In what follows, we will call $C_{n}$ a zigzag event. We refer to $n$ as the “center” of this event $C_{n}$. It is composed of the “zig” event $C_{n,1}$, of “length” $k$ (i.e. composed of $k$ events), and then of the “zag” event and $C_{n,2}$, also of length $k$. The length of the zigzag event is hence $2k$. Note that, for any $m$ and for any $1\leq i_{1},\ldots,i_{m}\leq k$ and $n_{1}<n_{2}<\ldots<n_{m}$, the events $(A_{i_{j},n_{j}})_{1\leq j\leq m}$ are mutually independent. Therefore, for any $n$, the events $C_{n,1}$ and $C_{n,2}$ are mutually independent and occur with the same (positive) probability.
Fig. 2 illustrates a sequence $\{B_{i}\}_{i=1}^{k}$ that forms a zigzag event for $\cal X\subset\mathbb{R}$ satisfying assumptions (a)-(c).
A zigzag event creates an isolation between customers who arrive before the zig and customers who arrive after the zag, where the latter are served only after the former have left the system. Based on zigzag events, we now introduce some splitting of customers into “blocks”, or “cycles”. Let $n_{0}=0$, $n_{1}=2k\cdot\min\{r\geq 1:{\mathbf{I}}(C_{2kr})=1\}$ and, for $m=1,2,\ldots$,
$n_{m+1}=n_{m}+2k\cdot\min\{r\geq 1:{\mathbf{I}}(C_{n_{m}+2kr})=1\},$ (8)
where $\mathbf{I}$} is the indicator function, which is equal to 1 if the event occurs and 0 otherwise. In words, $n_{m}$ is the center of the $m$-th zigzag event on the integer lattice of multiples of $2k$.
@@
  • The last customer of this block (which also has an exclusion set containing $B_{k}$) is the closing customer of block $m$; it arrives at $\widehat{T}_{m}:=t_{n_{m+1}-1}$.
For $m\geq 0$, let
$\beta_{m}:=n_{m+1}-n_{m},$
which denotes the number of customers in block $m$. The distribution of $\{\beta_{m}\}_{m\geq 0}$ is characterized by the lemma below.
Lemma 5.
$\beta_{0},\beta_{1},$... are i.i.d. and geometrically distributed random variables with parameter $p_{b}$ multiplied by $2k$, where
$p_{b}:={\mathbb{P}}(C_{n})=\left(\prod_{i=1}^{k}{\mathbb{P}}(B_{i}\subseteq S)\right)^{2}>0.$
Proof.
Recall that the exclusion sets $\{S_{i}\}$ are independent. The events $\{C_{2k},C_{4k},C_{6k},$...\} are stationary and mutually independent, since each of these events depends on exactly $2k$ exclusion sets with indices drawn from disjoint sets. Let $r_{1}=\min\{r\geq 1:{\mathbf{I}}(C_{2kr})=1\}$. $r_{1}$ is the index of the first success in Bernoulli trials ($C_{2kr_{1}}$ is the first zigzag event) with parameter $p_{b}$ and is geometrically distributed. The fact that $(n_{m+1}-n_{m})/(2k)=((n_{m+1}+k-1)-(n_{m}+k-1))/(2k)$ is i.i.d. and geometrically distributed follows from the definition in Eq (8) and a similar argument as that of $r_{1}$. ∎
Lemma 5 shows that the random variables $\beta_{0},\beta_{1},\ldots$ and $n_{1},n_{2},\ldots$ are all a.s. finite. Specifically, ${\mathbb{E}}\beta_{m}=2k/p_{b}.$
Let $T_{0}=0$ and $T_{m}:=t_{n_{m}}$. For $m\geq 0$, let
$Y_{m}:=\{((x_{n},t_{n+1}-t_{n_{m}}),(S_{n},h_{n})),n=n_{m},n_{m}+1,n_{m}+2,\ldots,n_{m+1}-1\}.$
(a) An illustration of a block of arrivals, $k=4$.
(b) An equivalent representation of block $m$ viewed at the beginning of its service when waiting time $W_{m}\geq\widehat{T}_{m}-T_{m}$.
Figure 3: An illustration of the block structure for $\cal X\subset\mathbb{R}$ and $k=4$, where there is one (non-zigzag) set of arrivals of size $2k$ between two zigzags and they are indexed by order of arrival. Arrivals with the same color have the same exclusion radius and file size (height). In the left figure, the base of a rectangle indicates the arrival time. In the right figure, for $W_{m}\geq\widehat{T}_{m}-T_{m}$, an equivalent representation of block $m$ is depicted. At the beginning of the service of the block, all arrivals are already present, hence the block is in effect maximally compressed in time while still preserving the local FCFS mechanism.
Here are a few observations about this definition:
  • By the definition in (8), the size of each block $Y_{m},m\geq 1$, is an integer multiple of $2k$.
  • The index of the $m+1$-st block is an integer larger than or equal to $n_{m}+2k$ (it is equal when there are two consecutive zigzag events, with their centers separated by $2k$).
  • A block begins with a “zag” sub-block of size/length $k$, followed by $\beta_{m}/(2k)-1$ “intermediate” sub-blocks (each of length $2k$, none of which being a zigzag event), and ends with a “zig” sub-block of length $k$. All these sub-blocks are mutually independent, with intermediate blocks being identically distributed. There is no intermediate sub-block if $\beta_{m}/(2k)=1$.
  • These blocks are mutually independent and identically distributed, except for $Y_{0}$. The block $Y_{0}$ differs from the others only in the first sub-block of size $k$, because its elements have the unconditional distribution.
The following proposition is a summary of what has to be retained from the last observations:
@@
The sequence of random elements $\{Y_{m}\}_{m\geq 1}$ is i.i.d. and independent of $Y_{0}$.
@@
There are several other options for defining zigzag events while preserving the properties identified above and below. One of them is that with no-doubling of the last event in the zig. For example, if $k=1$ (i.e., we have customers with an exclusion set equal to the whole space), then we do not have to double this event and all the properties identified still hold. This is the case when $\{R_{i}\}$ is i.i.d. with a distribution with unbounded support, e.g., exponential.
The specific structure proposed here is retained for its simplicity and symmetry. This concept comes from tropical mathematics and is linked to what is now called the tropical matrix rank. See, e.g., [12].
Note also that, in Subsection III-A, Assumption A3, we may add for free a further assumption (e) to assumptions (a)–(d):
@@
Indeed, if (e) fails for some  $i\neq j$, then we may merge these two sets and consider a new collection of (k1)(k-1) events $B_{i}$’s instead.
There are other options for defining the blocks. For example, we may deal with smaller blocks, by defining $\{n_{m}\}$ inductively as
@@
In this case, the number of arrivals in each block does not need to be a multiple of $2k$. However, we prefer the definition in (8) because the variant with (9) leads to a more complicated construction.
III-C Reduction to a Lindley-Type Recurrence Relation
In view of the service discipline and the definition of blocks, customers are served block by block, i.e., the service of the customers of block $m+1$ may start only when the service of the closing customer of block $m$ is completed and, at that time, all customers of block $m$ have already left. This can be exploited to derive a Lindley-type evolution equation.